Sometimes we want a confidence level using the T-Distribution. What is t a/2 when there are 34 degrees of freedom and the confidence level is 99%?

For this problem, round your answers to 3 digits after the decimal point
Answer #1

degrees of freedom, df = n - 1 = 34

Given CI level is 99%, hence α = 1 - 0.99 = 0.01
α/2 = 0.01/2 = 0.005, tc = t(α/2, df) = 2.728